OpenStudy (anonymous):

Ok so 48 squares are in the box and 14 squares of the 48 are colored red what percentage are colored red please help i know im young i need help!! I dont understand

OpenStudy (anonymous):

14/48 * x/100 so cross multiply 14*100 then divide by 48

OpenStudy (anonymous):

ok.. is that all??

jimthompson5910 (jim_thompson5910):

part/whole = percent/100 14/48 = x/100 14*100 = 48*x ... cross multiply 1400 = 48x 48x = 1400 x = ???

OpenStudy (anonymous):

29%

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

yes x = 29.1667 and that rounds to 29%

jimthompson5910 (jim_thompson5910):

yw